heloo guys i have these homework assignment i need help on it..

1- A car moving initially at 48 mi/h begins decelerating at a constant rate 110 ft short of a stoplight.If the car comes to a full stop just at the light, what is the magnitude of its acceleration?

Express your answer to two significant figures and include the appropriate units.

2- A gallon of paint covers 420 ft2 .What is its coverage in m2/L?

3- A car moving initially at 48 mi/h begins decelerating at a constant rate 110 ft short of a stoplight.

If the car comes to a full stop just at the light, what is the magnitude of its acceleration?

Express your answer to two significant figures and include the appropriate units.

4- A jetliner leaves San Francisco for New York, 4600 km away. With a strong tailwind, its speed is 1040 km/h . At the same time, a second jet leaves New York for San Francisco. Flying into the wind, it makes only 780 km/h .When do the two planes pass each other?Where do the two planes pass each other?

5- A police radar's effective range is 1.2 km , and your radar detector's range is 2.1 km . You're going 110 km/h in a 74 km/h zone when the radar detector beeps.

At what rate must you negatively accelerate to avoid a speeding ticket?

Express your answer to two significant figures and include the appropriate units.
Answer #1

Initial speed u = 48 mi/h

                     = 48(1609 m) /(3600 s)

                     = 21.453 m/s

Final speed v = 0

Stopping distance S = 110 ft

                              = 110(0.3048m)

                              = 33.528 m

From the relation v 2 - u 2 = 2aS

Acceleration a = [v 2 - u 2 ]/2S

                      = [0 2 - 21.453 2 ] /(2x33.528]

                      = - 6.863 m/s 2

                      = -6.9 m/s 2